# Efficient Conformal Prediction under Data Heterogeneity

### v1

@@ 1 Introduction
The equation (1) essentially bounds the miscoverage rate on average over possible sets of calibration data and test points. However, if there is a high variability in the coverage probability as a function of the calibration data, the test coverage probability may be substantially below $1-\alpha$ for a particular calibration set. Therefore, for a given calibration set $\mathcal{D}_{N}$ the reliability of the confidence set can be assessed via the empirical miscoverage rate:
@@ 2.1 Basics of Conformal Prediction
$\mathcal{C}_{\alpha,\mu}(\mathbf{x})\{\mathbf{y}\in\mathcal{Y}\colon V(\mathbf{x},\mathbf{y})\leq Q_{1-\alpha}(\mu)\},$ (6)
@@ 2.2 Efficient Conformal Prediction under Data Shift
Notably, using $\cosh(\theta q_{k})\leq\mathrm{e}^{(\theta q_{k})^{2}/2}$ with $q_{k}=1/N$ leads to the standard DKW result. However, utilizing $\cosh(\theta q_{k})\leq\exp(2^{-1}\theta^{2}\mathbb{E}[q_{k}^{2}])\cosh\{\theta(q_{k}-\mathbb{E}q_{k})\}$ provides a sharper result:
@@ 5 Experiments
In our testing phase, we set $\alpha=0.1$ and, for the non-conformity score, we adopt the adaptive prediction sets (APS) approach developed in (Romano et al.,, 2020; Angelopoulos et al.,, 2023): {EQA}[c] V(x, y)= ∑{y:^f(x)y^f(x)y} ^f(x)y where $\hat{f}(\mathbf{x})_{y}$ is the predicted probability of label $y\in\mathcal{Y}$. We consider ResNet-18 model trained on the unaltered CIFAR-10 dataset. We found temperature scaling of the logits beneficial and set the temperature $T=10$ in our experiments. Weights of the nonconformity scores are computed with the GMM approach described above. We compare our method against two baselines: “Unweighted Local” and “Unweighted Global”. Both use standard conformal prediction with $\lambda=({\text{Num data}+1})^{-1}$. The former only uses the local agent’s calibration data, while the latter combines calibration data from all agents. Our results are obtained from 100 independent runs, each using different random splits of an agent’s data for calibration, testing, and for density model fitting. For each iteration, we evaluate the local test set of an agent, calculating both average coverage and average set size. The metrics are collected across 100 runs and visualized using boxplots. Results for $\alpha=0.1$ are shown in Figure 1(a) and Figure 1(b). It is worth noting that while we had 40 agents, we aggregated results for clarity, presenting them by the level of data corruption. The figures show that except for “Unweighted Global”, all methods achieve the target coverage. However, our method consistently shows less variation in set sizes than the others. This consistency arises from the federated process, when calibration scores from other agents, once appropriately weighted, contribute to the conformal procedure for a given agent.
@@
We compare the performance of the introduced algorithm with local baselines at $5$ different corruption levels in terms of coverage and average prediction set size. Data allocation scheme is as follows. ImageNet and ImageNet-C validation data are distributed among $20$ agents without overlap, such that half of the agents contain clean, non-corrupted data, and the other half consists of corrupted images of the same level of corruption. Each agent contains $850$ data samples, of which $50$ are calibration samples, $300$ test samples and $500$ density model fitting samples. We conduct separate experiments with the above-mentioned setup for all levels of corruption. The results are presented as box plots with means and standard deviations obtained from $10$ random splits of the data for each of these experiments with varying levels of corruption. The results of the algorithms at each corruption level are obtained by averaging the results from all agents within the same level of corruption. The Figure 3(a) shows the empirical coverage of the prediction sets at different levels of corruption. The resulting coverage of the introduced method is valid in comparison with “Unweighted Global” baseline, has lower variance and better accuracy compare to the “Unweighted Local” one. The Figure 3(b) demonstrates the average prediction set size of conformal sets as a function of the level of corruption. The presented method shows on average slightly smaller set sizes and much less variance of the results compared to an “Unweighted Local” baseline. By using information from other agents through importance weights, our method overcomes the problems of data heterogeneity compared to unweighted alternatives.
@@ Appendix A Proof of the conditional coverage
$\mathbb{E}\left[{\exp\left({\theta\epsilon_{k}q_{k}}\right)}\right]$ $=\mathbb{E}\left[{\exp\left({\theta\epsilon_{k}\mathbb{E}{q_{k}}}\right)}\right]\mathbb{E}\left[{\exp\left({\theta\epsilon_{k}\left[{q_{k}-\mathbb{E}{q_{k}}}\right]}\right)}\right]$
$\leq\exp\left({2^{-1}\theta^{2}\mathbb{E}[{q_{k}}]^{2}}\right)\mathbb{E}\left[{\exp\left({\theta\epsilon_{k}\left[{q_{k}-\mathbb{E}{q_{k}}}\right]}\right)}\right].$
@@
$\prod_{k=1}^{N}\mathbb{E}\left[{\cosh\left({\theta q_{k}}\right)}\right]$ $\leq\exp\left({\frac{\theta^{2}}{2}\sum_{k=1}^{N}\mathbb{E}[{q_{k}}]^{2}}\right)\mathbb{E}\left[{\mathbb{E}\left[{\exp\left({\theta\sum_{k=1}^{N}\epsilon_{k}\left[{q_{k}-\mathbb{E}{q_{k}}}\right]}\right)\,\bigg{|}\,\left\{{\epsilon_{k}}\right\}_{k\in[N]}}\right]}\right]$
@@
$\prod_{k=1}^{N}\mathbb{E}\left[{\cosh\left({\theta\tilde{q}_{k}}\right)}\right]$ $\leq\exp\left({\frac{\theta^{2}}{2}\sum_{k=1}^{N}\mathbb{E}[{\tilde{q}_{k}}]^{2}}\right)\mathbb{E}\left[{\mathbb{E}\left[{\exp\left({\theta\sum_{k=1}^{N}\epsilon_{k}\left[{\tilde{q}_{k}-\mathbb{E}{\tilde{q}_{k}}}\right]}\right)\,\bigg{|}\,\left\{{\epsilon_{k}}\right\}_{k\in[N]}}\right]}\right]$
$\leq\exp\left({\frac{\theta^{2}}{2({\sum_{l=1}^{N+1}\mathbb{E}\lambda_{l}})^{2}}\sum_{k=1}^{N}\mathbb{E}[{\lambda_{k}}]^{2}+\frac{2\theta^{2}}{({\sum_{l=1}^{N+1}\mathbb{E}\lambda_{l}})^{2}}\sum_{k=1}^{N}\sigma_{k}^{2}}\right).$ (29)
@@ Appendix B Extension of the DKW theorem
$\mathbb{E}\left[{\exp\left({\theta\sup_{z\in\mathcal{Z}}\left\{{\hat{G}_{N}(z)-G(z)}\right\}}\right)}\right]\leq 2\exp\left({2\theta^{2}\sum_{k=1}^{N}(g_{k}(Z_{k}))^{2}}\right).$
Plugging the previous inequality into (34), and using the independence between $\{g_{k}(Z_{k})\}_{k=1}^{N}$ implies
$\mathbb{P}\left({\sup_{z\in\mathcal{Z}}\left\{{\hat{G}_{N}(z)-G(z)}\right\}}\right)\leq 2\inf_{\theta>0}\left\{{\mathrm{e}^{-\theta\epsilon}\prod_{k=1}^{N}\mathbb{E}\left[{\cosh\left({\theta g_{k}(Z_{k})}\right)}\right]}\right\}.$
@@
Let $\{\epsilon_{k}\}_{k\in[N]}$ be i.i.d. random Rademacher variables independent of $\{(Z_{k},\tilde{Z}_{k})\}_{k=1}$, we have
@@
$\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}=\begin{cases}0&\text{ if }V(z)<V(Z_{\sigma(1)})\\ \sum_{j=1}^{i}\epsilon_{\sigma(j)}p_{\sigma(j)}^{Z_{\sigma(j)}}&\text{ if }V(Z_{\sigma(i)})\leq V(z)<V(Z_{\sigma(i+1)})\\ \sum_{j=1}^{N}\epsilon_{\sigma(j)}p_{\sigma(j)}^{Z_{\sigma(j)}}&\text{ if }V(z)\geq V(Z_{\sigma(N)})\end{cases}.$
@@
$\sup_{z\in\mathcal{Z}}\exp\left({2\theta\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}}\right)\leq\sup_{0\leq i\leq n}\exp\left({2\theta\sum_{j=1}^{i}\epsilon_{\sigma(j)}p_{\sigma(j)}^{Z_{\sigma(j)}}}\right).$
@@
$\mathbb{E}\left[{\sup_{z\in\mathcal{Z}}\exp\left({2\theta\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}}\right)\,\bigg{|}\,\{{Z_{k}}\}_{k=1}^{N}}\right]\\ \leq\mathbb{E}\left[{\sup_{0\leq i\leq N}\exp\left({2\theta\sum_{j=1}^{i}\epsilon_{\sigma(j)}p_{\sigma(j)}^{Z_{\sigma(j)}}}\right)\,\bigg{|}\,\{{Z_{k}}\}_{k=1}^{N}}\right]\leq 2\prod_{k=1}^{N}\cosh\left({\theta g_{k}(Z_{k})}\right).$
@@
$=\frac{\sum_{k=1}^{N}\{{\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}-\mathbb{E}[{\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}\,|\,V_{N+1}}]}\}}{\sum_{l=1}^{N+1}\lambda({Z_{l}})}+\frac{\sum_{l=1}^{N+1}\{{\lambda({Z_{l}})-\lambda({\tilde{Z}_{l}})}\}}{\left({\sum_{l=1}^{N+1}\lambda({Z_{l}})}\right)\left({\sum_{l=1}^{N+1}\lambda({\tilde{Z}_{l}})}\right)}\mathbb{E}\left[{\sum_{k=1}^{N}\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}\,\bigg{|}\,V_{N+1}}\right]$
@@
[…]

### v2

@@ 1 Introduction
Here the notation $\mathbb{P}$ denotes that the probability is computed with respect to $\{(X_{i},Y_{i})\}_{i=1}^{N}$ and $(X_{N+1},Y_{N+1})$, which are assumed i.i.d. with common distribution $P$. The equation (1) essentially bounds the miscoverage rate on average over possible sets of calibration data and test points. However, if there is a high variability in the coverage probability as a function of the calibration data, the test coverage probability may be substantially below $1-\alpha$ for a particular calibration set. Therefore, for a given calibration set $\mathcal{D}_{N}$ the reliability of the confidence set can be assessed via the empirical miscoverage rate:
@@ 2.1 Basics of Conformal Prediction
$\mathcal{C}_{\alpha,\mu}(\mathbf{x})=\{\mathbf{y}\in\mathcal{Y}\colon V(\mathbf{x},\mathbf{y})\leq Q_{1-\alpha}(\mu)\},$ (6)
@@ 2.2 Efficient Conformal Prediction under Data Shift
Notably, using $\cosh(\theta q_{k})\leq\mathrm{e}^{(\theta q_{k})^{2}/2}$ with $q_{k}=1/N$ leads to the standard DKW result. However, utilizing $\mathbb{E}\cosh(\theta q_{k})\leq\cosh(\theta\mathbb{E}q_{k})\exp(2\theta^{2}\sigma_{k}^{2}/N^{2})$ provides the following result:
@@ 5 Experiments
In our testing phase, we set $\alpha=0.1$ and, for the non-conformity score, we adopt the adaptive prediction sets (APS) approach developed in (Romano et al.,, 2020; Angelopoulos et al.,, 2023): {EQA}[c] V(x, y)= ∑{y:^f(x)y^f(x)y} ^f(x)y, where $\hat{f}(\mathbf{x})_{y}$ is the predicted probability of label $y\in\mathcal{Y}$. We consider ResNet-18 model trained on the unaltered CIFAR-10 dataset. We found temperature scaling of the logits beneficial and set the temperature $T=10$ in our experiments. The weights associated to the nonconformity scores are computed with the GMM approach described above. We compare our method against three baselines: “Unweighted Local”, “Unweighted Global” and “Weighted global (resampled)” (Plassier et al.,, 2023). First two use standard conformal prediction with $\lambda=({\text{Num data}+1})^{-1}$, while the latter uses specially constructed weights. “Unweighted Local” only uses the local agent’s calibration data, while “Unweighted Global” combines calibration data from all agents. Our results are obtained from 100 independent runs, each using different random splits of an agent’s data for calibration, testing, and for density model fitting. For each iteration, we evaluate the local test set of an agent, calculating both average coverage and average set size. The metrics are collected across 100 runs and visualized using boxplots.
Results for $\alpha=0.1$ are shown in Figure 2(a) and Figure 2(b). It is worth noting that while we had 40 agents, we aggregated results for clarity, presenting them by the level of data corruption. The figures show that except for “Unweighted Global”, all methods achieve the target coverage. “Weighted global (resampled)” has a coverage of $92.28\pm 5.97\%$ which is perfectly aligned with expectation to have increased variance compared to our method (see Table 1). In contrast, our method consistently shows less variation in set sizes than the others. This consistency arises from the federated process, when calibration scores from other agents, once appropriately weighted, contribute to the conformal procedure for a given agent.
It is worth mentioning, that we do not present (Tibshirani et al.,, 2019) here among competitors. For this method, the only feasible approach is to sample a certain number of permutations and compute approximate weights based on them. Unfortunately, it leads to the very high variance of the weights which led us even to the bias in the mean coverage ($82.35$ on the “Mix” calibration data in our experiment on domain adaptation). When replicating the CIFAR10 experiment, sampling the permutations led us to even worse results.
@@
We compare the performance of the introduced algorithm with local baselines at $5$ different corruption levels in terms of coverage and average prediction set size. Data allocation scheme is as follows. ImageNet and ImageNet-C validation data are distributed among $20$ agents without overlap, such that half of the agents contain clean, non-corrupted data, and the other half consists of corrupted images of the same level of corruption. Each agent contains $850$ data samples, of which $50$ are calibration samples, $300$ test samples and $500$ density model fitting samples. We conduct separate experiments with the above-mentioned setup for all levels of corruption. The results are presented as box plots with means and standard deviations obtained from $10$ random splits of the data for each of these experiments with varying levels of corruption. The results of the algorithms at each corruption level are obtained by averaging the results from all agents within the same level of corruption.
The Figure 4(a) shows the empirical coverage of the prediction sets at different levels of corruption. The resulting coverage of the introduced method is valid in comparison with “Unweighted Global” baseline, has lower variance and better accuracy compare to the “Unweighted Local” one. The Figure 4(b) demonstrates the average prediction set size of conformal sets as a function of the level of corruption. The presented method shows on average slightly smaller set sizes and much less variance of the results compared to an “Unweighted Local” baseline. By using information from other agents through importance weights, our method overcomes the problems of data heterogeneity compared to unweighted alternatives.
@@ 6 Conclusion
Acknowledgements
The authors gratefully thank Pierre Humbert and Sylvain Arlot for their valuable feedback. The research on experimental evaluation of proposed methods (Section 5) was supported by the Russian Science Foundation grant 20-71-10135. Part of this work has been carried out under the auspice of the Lagrange Mathematics and Computing Research Center. The work of F. Noskov was prepared within the framework of HSE basic research program.
@@ Appendix A Proof of the conditional coverage
$\mathbb{E}\left[{\exp\left({\theta\epsilon_{k}\mathbb{E}{q_{k}}}\right)}\right]\leq\exp\left({2^{-1}\theta^{2}\mathbb{E}[{q_{k}}]^{2}}\right).$
@@
$\prod_{k=1}^{N}\mathbb{E}\left[{\cosh\left({\theta q_{k}}\right)}\right]$ $=\mathbb{E}\left[{\exp\left({\theta\sum_{k=1}^{N}\epsilon_{k}\mathbb{E}q_{k}}\right)\mathbb{E}\left[{\exp\left({\theta\sum_{k=1}^{N}\epsilon_{k}\left[{q_{k}-\mathbb{E}{q_{k}}}\right]}\right)\,\bigg{|}\,\left\{{\epsilon_{k}}\right\}_{k\in[N]}}\right]}\right]$
$\leq\mathbb{E}\left[{\exp\left({\theta\sum_{k=1}^{N}\epsilon_{k}\mathbb{E}q_{k}}\right)}\right]\exp\left({\frac{2\theta^{2}}{({\sum_{l=1}^{N}\mathbb{E}\lambda_{l}})^{2}}\sum_{k=1}^{N}\sigma_{k}^{2}}\right)$
@@
$\prod_{k=1}^{N}\mathbb{E}\left[{\cosh\left({\theta\tilde{q}_{k}}\right)}\right]\leq\exp\left({\frac{\theta^{2}}{2({\sum_{l=1}^{N+1}\mathbb{E}\lambda_{l}})^{2}}\sum_{k=1}^{N}\mathbb{E}[{\lambda_{k}}]^{2}+\frac{2\theta^{2}}{({\sum_{l=1}^{N+1}\mathbb{E}\lambda_{l}})^{2}}\sum_{k=1}^{N}\sigma_{k}^{2}}\right).$ (29)
@@ Appendix B Extension of the DKW theorem
Plugging the previous inequality into (34), and minimizing the resulting expression with respect to $\theta$ yields:
@@
Let $\{\epsilon_{k}\}_{k\in[N]}$ be i.i.d. random Rademacher variables independent of $\{(Z_{k},\tilde{Z}_{k})\}_{k=1}^{N}$, we have
@@
$\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}=\begin{cases}0&\text{ if }V(z)<V(Z_{\sigma(1)})\\ \sum_{j=1}^{i}\epsilon_{\sigma(j)}g_{\sigma(j)}(Z_{(j)})&\text{ if }V(Z_{\sigma(i)})\leq V(z)<V(Z_{\sigma(i+1)})\\ \sum_{j=1}^{N}\epsilon_{\sigma(j)}g_{\sigma(j)}(Z_{(j)})&\text{ if }V(z)\geq V(Z_{\sigma(N)})\end{cases}.$
@@
$\sup_{z\in\mathcal{Z}}\exp\left({2\theta\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}}\right)\leq\sup_{0\leq i\leq n}\exp\left({2\theta\sum_{j=1}^{i}\epsilon_{\sigma(j)}g_{\sigma(j)}(Z_{(j)})}\right).$
@@
$\mathbb{E}\left[{\sup_{z\in\mathcal{Z}}\exp\left({2\theta\sum_{k=1}^{N}\epsilon_{k}g_{k}(Z_{k})\mathds{1}_{V(Z_{k})\leq V(z)}}\right)\,\bigg{|}\,\{{Z_{k}}\}_{k=1}^{N}}\right]\\ \leq\mathbb{E}\left[{\sup_{0\leq i\leq N}\exp\left({2\theta\sum_{j=1}^{i}\epsilon_{\sigma(j)}g_{\sigma(j)}(Z_{(j)})}\right)\,\bigg{|}\,\{{Z_{k}}\}_{k=1}^{N}}\right]\leq 2\prod_{k=1}^{N}\cosh\left({\theta g_{k}(Z_{k})}\right).$
@@
$=\frac{\sum_{k=1}^{N}\{{\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}-\mathbb{E}[{\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}\,|\,V_{N+1}}]}\}}{\sum_{l=1}^{N+1}\lambda({Z_{l}})}$
$\qquad+\frac{\sum_{l=1}^{N+1}\{{\lambda({Z_{l}})-\lambda({\tilde{Z}_{l}})}\}}{\left({\sum_{l=1}^{N+1}\lambda({Z_{l}})}\right)\left({\sum_{l=1}^{N+1}\lambda({\tilde{Z}_{l}})}\right)}\mathbb{E}\left[{\sum_{k=1}^{N}\lambda({Z_{k}})\mathds{1}_{V_{k}<V_{N+1}}\,\bigg{|}\,V_{N+1}}\right]$
@@
$\left\lvert{\mathbb{P}\left({V_{N+1}\leq Q_{1-\alpha}\left({\textstyle\sum_{k=1}^{N}p_{k}\delta_{V_{k}}+p_{N+1}\delta_{\infty}}\right)}\right)-1+\alpha}\right\rvert\leq\begin{cases}35\left\|{\lambda}\right\|_{\infty}^{2}\sqrt{\frac{\log 4N}{N}}&\text{if $\lambda$ is bounded}\\ \frac{18\mathbb{E}\lambda^{2}(Z_{N+1})}{N}+19\sigma\sqrt{\frac{\log 4N}{N}}&\text{otherwise}\end{cases}.$
@@ E.3 Additional experiment on CIFAR-100
E.4 Comparison with Tibshirani’s weights
Tibshirani et al., (2019)’s algorithm relies on importance weights derived from the density of $(Z_{\sigma(1)},\ldots,Z_{\sigma(N+1)})$ calculated over all possible permutations $\sigma$, that in general case requires the evaluation of sums of $N!$ terms, where $N$ represents the total number of calibration points. Tibshirani’s expression simplifies in very special cases, such as when the distribution of the test point is different from the distribution of the calibration data, which are i.i.d. (an example discussed in Tibshirani et al., (2019)). Our proposed conformal prediction set aligns with Tibshirani’s when the calibration data are i.i.d.. However, our main contribution is that our prediction set remains valid even for the non-i.i.d. calibration data, i.e. in the presence of label and covariate shifts inside the calibration data. To show this, we provide conditional coverage guarantees and avoid the intractability of combinatorial Tibshirani’s CP set.
Plassier et al., (2023) method consists in subsampling the calibration data to generate i.i.d. samples from the mixture distribution, and then using the “simplified” Tibshirani’s weights. However, the subsampling cost leads to an increase in variance (and also the theory becomes more convoluted compared to our paper).
We conducted additional experiments comparing the work of Plassier et al., (2023) and Tibshirani et al., (2019). For Tibshirani’s method, the only feasible approach is to sample a certain number of permutations and compute approximate weights based on them. Unfortunately, it leads to the very high variance of the weights which led us even to the bias in the mean coverage ($82.48$ on the “Mix” calibration data in our experiment on domain adaptation). For Plassier’s method we observed a coverage of $92.28\pm 5.97\%$ which is perfectly aligned with expectation to have increased variance compared to our method; see Figure 1. When replicating the CIFAR10 experiment, sampling the permutations led us to even worse results and we only present the comparison with Plassier’s method which again gives an expected increase in variance (see figures below).
Figure 8: Coverage comparison on CIFAR-10. (Left) Empirical coverage in function of the data corruption level. (Right) Average set size of conformal prediction sets as a function of data corruption level.